# Polaronic dressing of bound states

### v1

@@ Abstract
Introduction.- The polaron remains a powerful tool to understand the behavior of isolated impurities in complex many-body systems Landau and Pekar (1948). The versatility of this concept has allowed the polaron to be exploited in several fields in physics spanning from atomic physics Jørgensen et al. (2016); Hu et al. (2016), solid-state Pekar (1969); Eagles (1963); Devreese and Alexandrov (2009), quantum optics Grusdt and Fleischhauer (2016); Nielsen et al. (2020); Camacho-Guardian et al. (2020), chemistry Srimath Kandada and Silva (2020); Franchini et al. (2021); Tao et al. (2021); Bourelle et al. (2022) to nuclear physics Kobyakov and Pethick (2016); Tajima et al. (2024).
@@
Ultracold quantum gases provide a powerful and reliable platform to quantum simulate complex many-body phenomena due to the high controllability of interactions. Thus, polarons emerging from the interaction of impurities with bosonic quantum gases have significantly increased our understanding and allowed us to explore previously inaccessible regimes in solid-state setups, ruled by strong correlations Jørgensen et al. (2016); Hu et al. (2016); Ardila et al. (2019); Yan et al. (2020); Skou et al. (2021a, 2022); Etrych et al. (2024). This opened up a plethora of new intriguing phenomena which most of them remain to be fully elucidated, for instance, the underlying character of the strongly interacting polaron Rath and Schmidt (2013); Ardila and Giorgini (2015); Shchadilova et al. (2016); Grusdt et al. (2018); Christensen et al. (2015); Christianen et al. (2023); Levinsen et al. (2021), thermal fluctuations effects Levinsen et al. (2017); Guenther et al. (2018); Field et al. (2020); Drescher et al. (2021); Isaule (2024), dipolar and charged polarons Ardila and Pohl (2018); Ardila (2022); Astrakharchik et al. (2021); Christensen et al. (2021); Ding et al. (2022); Astrakharchik et al. (2022); Cavazos Olivas et al. (2024); Volosniev et al. (2023), Efimov physics Levinsen et al. (2015); Sun et al. (2017); Sun and Cui (2017); Christianen et al. (2022a, b), mediated interactions Naidon (2018); Dehkharghani et al. (2018); Camacho-Guardian et al. (2018); Camacho-Guardian and Bruun (2018); Mistakidis et al. (2020); Keiler et al. (2021); Paredes et al. (2024); Levinsen et al. (2024), universality Yoshida et al. (2018); Massignan et al. (2021); Yegovtsev et al. (2022); Skou et al. (2022, 2021b); Cayla et al. (2023), lattice polarons Colussi et al. (2023); Ding et al. (2023); Isaule et al. (2024); Amelio et al. (2024); Santiago-García and Camacho-Guardian (2023); Santiago-García et al. (2024); Pimenov (2024); Vashisht et al. (2024); Amelio et al. (2024); Pimenov et al. (2021); Christ et al. (2024), among others Nielsen and Ardila (2019); Guenther et al. (2021); big (2022); Camacho-Guardian (2023); Mostaan et al. (2023). Recognizing the coupling of internal degrees of freedom to a quantum reservoir led to the identification of a new class of quasiparticle termed angulon Schmidt and Lemeshko (2015, 2016).
In many realistic cases, the impurity might possess a structure and it is formed by more than one elementary block. A canonical example in semiconductors is the exciton: a compound object created by the binding of an electron with a hole. Electrons and holes can dress separately with phonons of the lattice forming an electron-polaron and hole-polaron quasiparticles, respectively; however, when combined, the resulting exciton may possess very different properties than the ones associated with the bare components ( electrons and holes). In this context, understanding the polaron effects on the formation of excitons attracted much attention, mainly in the context of the Fröhlich polarons Emin (2013). On the other hand, recent inorganic and hybrid semiconductors have challenged this picture as the polaronic effects may extend beyond the Fröhlich regime Thouin et al. (2019).
@@
Despite the increasing interest on the Bose polaron in the strongly interacting regime, much less is known about the effects of polaronic features on underlying few- and many - body states. The simplest and perhaps of most immediate interest is the study of the polaron effects emerging from a molecular state formed by two impurity atoms. Intrigued by this question, in this Article we study the features of a molecular arising from the binding of two atoms strongly coupled to a Bose-Einstein condensate (BEC). We address the two-body problem by solving the Bethe-Salpeter for the interaction of two impurity atoms. The impurity atoms strongly coupled to the BEC forming a polaron treated at the level of the non-self consistent T-matrix approximation. Our results demonstrate intricate two-body scattering properties in which the molecular picture can be broken, leading to a phase-diagram (see Fig. 1) where two-body states compete with polaron physics. The study of the dressing of two-body molecular states beyond the Fröhlich model may pave the way for further studies in organic and inorganic semiconductors.
@@
here $\hat{b}^{\dagger}_{\mathbf{k}}$ and $\hat{c}^{\dagger}_{\mathbf{k}}$ create a boson/impurity with momentum $\mathbf{k}$ and energy $\varepsilon^{(b/c)}_{\mathbf{k}}=\frac{k^{2}}{2m_{b/c}}.$ The impurity-boson interaction is assumed to be short-ranged and well characterized by coupling strength $g_{IB}=\frac{2\pi a_{IB}}{m_{r}}$, depending on the scattering length $a_{IB}$ and the reduced mass $m_{r}^{-1}=m_{b}^{-1}+m_{c}^{-1}$. The impurity-impurity interaction is also considered to be a contact interaction with $g_{II}=\frac{4\pi a_{II}}{m_{c}}$. We take the volume of the system $V=1.$
We focus on the polaron effects on the dimer state. In particular, in the absence of impurity-boson interaction, the impurity-impurity interaction supports a two-body bound state for $a_{II}>0$ with a binding energy of $\varepsilon_{I}^{(0)}=-1/m_{c}a_{II}^{2}$. To include the polaronic dressing of this bound state we study the Bethe-Salpeter equation for the impurity-impurity scattering, which is given by Fetter and Walecka (1971),
@@
Here, the spectral function is defined as $S(\mathbf{k},\omega)=-2\text{Im}\mathcal{G}_{cc}(\mathbf{k},\omega).$ In this formalism, the poles of the Bethe-Salpeter give the two-body bound states. To include the strong coupling between the impurity and the bath we take the Dyson’s equation for the impurity
$G^{-1}_{cc}(\mathbf{k},\omega)=\omega-\frac{k^{2}}{2m_{c}}-\Sigma_{cc}(\mathbf{k},\omega),$ (4)
and employ the NSCT approximation to start understanding this problem, we then take the analytical expression for the self-energy Rath and Schmidt (2013)
@@
here, the total mass is denoted by $M=m_{b}+m_{c}$. Under this theoretical approach, the polaron properties can be calculated as following: the quasiparticle energy $\omega_{\mathbf{k}}^{P}$ and residue $Z_{\mathbf{k}}^{P}$ are defined in terms of the poles $\text{Re}[\mathcal{G}^{-1}_{cc}(\mathbf{k},\omega_{\mathbf{k}}^{P})]=0$ and $Z_{\mathbf{k}}^{P}=[\partial_{\omega}\text{Re}[\mathcal{G}^{-1}_{cc}(\mathbf{k},\omega)]]_{\omega=\omega_{\mathbf{k}}^{P}}^{-1}$ respectively. Further details of the single polaron can be found in a recent review Grusdt et al. (2024).
The NSCT considers an ideal BEC, and while it accounts for the impurity-boson Feshbach resonance, it is restricted to one Bogoliubov excitation out of the condensate. This formalism Rath and Schmidt (2013), has turned to describe extremely well the first generation of the Bose polaron experiments  Jørgensen et al. (2016); Hu et al. (2016); Ardila et al. (2019); Yan et al. (2020); Skou et al. (2021a). Here, we take a vanishing boson-boson interaction $a_{BB}=0,$ for a non-zero boson-boson interaction the qualitative and quantitative quasiparticle properties of the polaron remain very similar in the NSCT approximation Rath and Schmidt (2013), and thus, our analysis holds for $a_{BB}>0.$
In the absence of polaron dressing, the Bethe-Salpeter acquires an analytical expression
@@
where the last expression only holds around the pole of the Bethe-Salpeter equation, which of course, lies at the energy of the bare dimer $\varepsilon_{I}^{(0)}=-1/m_{c}a_{II}^{2}$. Here, we will fix $m_{c}=m_{b}=m.$
@@
This intuitive picture is, however, not entirely accurate, and the relation between polaron physics and the underlying two-body physics is more profound, as we now demonstrate. For this purpose, we start by solving the full Bethe-Salpeter equation for a fixed value of $a_{II}$ and for varying impurity-boson interactions $a_{IB}.$ To understand the polaron effects on the scattering properties we introduce the spectral function of the scattering matrix defined as $A(\mathbf{Q},\omega)=-2\text{Im}\Gamma(\mathbf{Q},\omega)$ and take first $\mathbf{Q}=0.$
@@
We first study the case of $0.51/k_{n}a_{II}=$ 0.5, which corresponds to a binding energy of $\varepsilon_{I}^{(0)}/E_{n}=-0.5$. The spectral function of the impurity-impurity scattering matrix is shown in Fig. 2 as a function of $1/k_{n}a_{IB}$, which illustrates that the dressed dimer evolves from a well- to an ill-defined state. For weak boson-impurity interactions the spectral function is a sharp function with most of the spectral weight concentrated around the pole, giving a well-defined dressed dimer. As the polaron effects become more pronounced, we observe that the pole broadens significantly, hindering any signature of a dressed dimer state. This indicates that two-body bound states are well-defined when polaronic effects are weaker, whereas as polaron dressing increases, it prevents the emergence of well-defined dimers.
@@
In our work we have neglected induced interaction between the impurities, which however, can be strong enough to form bipolarons Naidon (2018); Camacho-Guardian et al. (2018); Camacho-Guardian and Bruun (2018) . We have assumed that the dynamics involved in the formation of the dimer due to direct interactions is much faster than the dynamics associated to mediated interactions. This approximation holds as long as the typical velocities associated to the exchange of phonons, namely, the speed of sound of the BEC, $c_{s}=\sqrt{4\pi a_{BB}n_{0}/m_{b}^{2}}$ remains much smaller than the typical velocities of the dimer, that is, $1\gg c_{s}/\sqrt{2|\varepsilon_{I}^{(0)}|/m_{b}}\sim\sqrt{n_{0}a_{BB}a_{IB}^{2}}.$ Here, $a_{BB}$ is the scattering length for the boson-boson interaction. That is, dimers with large binding energies are also less sensible to mediated interactions. In our case, we have studied the case of $a_{BB}=0,$ consistent with this discussion. We should stress that, contrary to the mediated interaction which is strongly affected by retardation effects, the two-body physics that gives rise to the formation of the dimer is instantaneous.
Our analysis reveals then an interplay between the underlying attractive and repulsive polaron branches which can strongly modify the dressed dimer state. Furthermore, the character of the bare dimer state determines how sensible is the bound state to the polaronic dressing.
@@
Conclusions.- We have studied the polaron effects on the formation of a bound state due to the direct interaction between two impurities. By using a Green’s function formalism which allowed us to the treat the single polaron non-perturbatively and by addressing the two-body problem by solving the Bethe-Salpeter for the interaction of two impurity atoms we have demonstrated a subtle interplay between polaron physics and an underlying two-body bound state.
@@
Our results show that for weakly bound dimers, polaron effects can be strong enough to completely break the formation of a dimer. On the other hand, tightly bound dimers are more robust towards polaron effects and remain well-defined even deep in the strongly interacting impurity-boson regime. This highlights the underlying composition of the two-body bound states. Our theoretical proposal can be tested with the current state-of-the-art experimental capabilities to probe Feshbach molecules Chin et al. (2010). An interesting avenue is to explore beyond the NSCT approximation the competition between polaron-few-body states Astrakharchik et al. (2021) and the formation of dressed dimers.
The study of the impact of beyond Fröhlich polaron dressing on two-body molecular state is relevant to other platform such as solid-state system or organic semiconductors where polaron features may dramatically change the properties of excitons, trions, and other few-body states. An interesting open question is to study this problem in the regime where the polaron description here employed breakdown.

### v3

@@ Abstract
Introduction.- The polaron remains a powerful tool to understand the behavior of isolated impurities in complex many-body systems Landau and Pekar (1948). The versatility of this concept has allowed the polaron to be exploited in several fields in physics spanning from atomic physics Jørgensen et al. (2016); Hu et al. (2016), solid-state Pekar (1969); Eagles (1963); Devreese and Alexandrov (2009), quantum optics Grusdt and Fleischhauer (2016); Nielsen et al. (2020); Camacho-Guardian et al. (2020), chemistry Srimath Kandada and Silva (2020); Franchini et al. (2021); Tao et al. (2021); Bourelle et al. (2022) to nuclear physics Kobyakov and Pethick (2016); Tajima et al. (2024). The arrival of ultracold quantum gases provided both a platform to quantum simulate complex many-body systems and to explore inaccessible regimes in solid-state setups, ruled by strong correlations Jørgensen et al. (2016); Hu et al. (2016); Ardila et al. (2019); Yan et al. (2020); Skou et al. (2021a, 2022); Etrych et al. (2024); Rath and Schmidt (2013); Ardila and Giorgini (2015); Shchadilova et al. (2016); Grusdt et al. (2018); Christensen et al. (2015); Christianen et al. (2023); Levinsen et al. (2021), thermal fluctuations effects Levinsen et al. (2017); Guenther et al. (2018); Field et al. (2020); Drescher et al. (2021); Hryhorchak and Pastukhov (2023); Isaule (2024), dipolar and charged polarons Ardila and Pohl (2018); Ardila (2022); Astrakharchik et al. (2021); Christensen et al. (2021); Ding et al. (2022); Astrakharchik et al. (2022); Cavazos Olivas et al. (2024); Volosniev et al. (2023), Efimov physics Levinsen et al. (2015); Sun et al. (2017); Sun and Cui (2017); Christianen et al. (2022a, b), universality Yoshida et al. (2018); Massignan et al. (2021); Yegovtsev et al. (2022); Skou et al. (2022, 2021b); Cayla et al. (2023), lattice polarons Colussi et al. (2023); Ding et al. (2023); Isaule et al. (2024); Amelio et al. (2024); Santiago-García and Camacho-Guardian (2023); Santiago-García et al. (2024); Pimenov (2024); Vashisht et al. (2024); Amelio et al. (2024); Pimenov et al. (2021); Christ et al. (2024), angulons Schmidt and Lemeshko (2015, 2016), among others Nielsen and Ardila (2019); Guenther et al. (2021); big (2022); Camacho-Guardian (2023); Mostaan et al. (2023).
In contrast to the remarkable theoretical and experimental progress in the single-impurity regime, much less is known about systems with two or more impurities. Considerable attention has been given to the theoretical study of mediated interactions Paredes et al. (2024), which challenge previously well-established formalisms from condensed matter physics Bardeen et al. (1967); Yu and Pethick (2012). This has led to the prediction of new forms of mediated interactions Camacho-Guardian and Bruun (2018); Fujii et al. (2022); Drescher et al. (2023) and strongly interacting bipolarons arising from the exchange of collective modes Naidon (2018); Camacho-Guardian et al. (2018); Will et al. (2021); Petković and Ristivojevic (2022); Drescher et al. (2023). Despite experimental efforts, clear evidence of mediated interactions and bipolarons has yet to be observed in the context of the Bose polaron, with only the first experiments signatures for Fermi polarons Baroni et al. (2024).
Bound states between two impurities are not limited to those formed via the exchange of sound modes in the medium; they may also arise from direct interactions between the impurities. A prime example is the exciton, a bound state between an electron and a hole in a semiconductor, which forms due to direct Coulomb attraction. Electrons and holes can individually couple to lattice phonons, forming electron-polaron and hole-polaron quasiparticles, respectively. However, when combined, the resulting exciton may exhibit properties significantly different from those of its bare constituents (electrons and holes). In this context, understanding the effects of polaron formation on exciton properties has drawn considerable attention, particularly in the framework of Fröhlich polarons Emin (2013).
@@
To the best of our knowledge, the effects of polaron dressing on existing bound states remain largely unexplored in the strongly interacting regime. Intrigued by this question, in this Article we study the features of a molecular arising from the binding of two atoms strongly coupled to a Bose-Einstein condensate (BEC). We address the two-body problem by solving the Bethe-Salpeter for the interaction of two impurity atoms. The impurity atoms strongly coupled to the BEC forming a polaron treated at the level of the non-self consistent T-matrix approximation. We reveal a phase diagram unveiling an unexplored competition between two-body and polaron physics.
@@
here $\hat{b}^{\dagger}_{\mathbf{k}}$ and $\hat{c}^{\dagger}_{\mathbf{k}}$ create a boson/impurity with momentum $\mathbf{k}$ and energy $\varepsilon^{(b/c)}_{\mathbf{k}}=\frac{k^{2}}{2m_{b/c}}.$ The impurity-boson interaction is assumed to be short-ranged and well characterized by coupling strength $g_{IB}=\frac{2\pi a_{IB}}{m_{r}}$, depending on the scattering length $a_{IB}$ and the reduced mass $m_{r}^{-1}=m_{b}^{-1}+m_{c}^{-1}$. The impurity-impurity interaction is also considered to be a contact interaction with $g_{II}=\frac{4\pi a_{II}}{m_{c}}$. Since we consider a s-wave interaction between the impurities we assume a bosonic statistic. We take the volume of the system $V=1.$
The direct interaction between impurities can be studied via the Bethe-Salpeter equation (BSE), which gives an exact solution for the two-body problem Fetter and Walecka (1971). In the absence of impurity-boson interaction, the impurity-impurity interaction supports a two-body bound state for $a_{II}>0$ with a binding energy of $\varepsilon_{I}^{(0)}=-1/m_{c}a_{II}^{2}$. The BSE is given by
@@
Here, the spectral function is defined as $S(\mathbf{k},\omega)=-2\text{Im}\mathcal{G}_{cc}(\mathbf{k},\omega).$ In this formalism, the poles of the BSE give the two-body bound states.
It is important to emphasize that while the BSE provides, in principle, the exact solution for impurity-impurity scattering, it depends on the impurity’s Green’s function, which can only be obtained approximately in the strong-coupling regime between the impurity and the bath. The impurity Green’s function follows Dyson’s equation
$G^{-1}_{cc}(\mathbf{k},\omega)=\omega-\frac{k^{2}}{2m_{c}}-\Sigma_{cc}(\mathbf{k},\omega).$ (4)
The problem of the strongly interacting Bose polaron remains an open question. In this article, we focus on understanding the effects of polaron dressing on the formation of two-body bound states. To this end, we employ a well-established theoretical formalism that accurately incorporates Feshbach physics Rath and Schmidt (2013), namely the so-called non-self-consistent T-matrix (NSCT) approximation.
The analytical expression for the self-energy Rath and Schmidt (2013)
@@
Here, we take a vanishing boson-boson interaction $a_{BB}=0,$ for a non-zero boson-boson interaction the qualitative and quantitative quasiparticle properties of the polaron remain very similar in the NSCT approximation Rath and Schmidt (2013), and thus, our analysis holds for $a_{BB}>0.$ Here, the total mass is denoted by $M=m_{b}+m_{c}$. Under this theoretical approach, the polaron properties can be calculated as following: the quasiparticle energy $\omega_{\mathbf{k}}^{P}$ and residue $Z_{\mathbf{k}}^{P}$ are defined in terms of the poles $\text{Re}[\mathcal{G}^{-1}_{cc}(\mathbf{k},\omega_{\mathbf{k}}^{P})]=0$ and $Z_{\mathbf{k}}^{P}=[\partial_{\omega}\text{Re}[\mathcal{G}^{-1}_{cc}(\mathbf{k},\omega)]]_{\omega=\omega_{\mathbf{k}}^{P}}^{-1}$ respectively. Further details of the single polaron can be found in a recent review Grusdt et al. (2024).
In the absence of polaron dressing, the BSE acquires an analytical expression
@@
where the last expression only holds around the pole of the BSE, which of course, lies at the energy of the bare dimer $\varepsilon_{I}^{(0)}=-1/m_{c}a_{II}^{2}$. Here, we will fix $m_{c}=m_{b}=m.$
@@
The BSE has been used to study bipolarons arising from the exchange of sound modes Camacho-Guardian et al. (2018). Due to the inherent complexity of the mediated interaction, which can be non-local and retarded, the BSE has so far been solved only within the quasiparticle approximation for the impurity’s Green’s function Camacho-Guardian et al. (2018). In this work, we go beyond this approximation by retaining the full Green’s function to solve the BSE, allowing us not only to determine the binding energy but also to address the full scattering problem.
To understand the polaron effects on the scattering properties we introduce the spectral function of the scattering matrix defined as $A(\mathbf{Q},\omega)=-2\text{Im}\Gamma(\mathbf{Q},\omega)$ and take first $\mathbf{Q}=0,$ and solve the BSE for a fixed value of $a_{II}$ and for varying impurity-boson interactions $a_{IB}.$
@@
We first study the case of $0.51/k_{n}a_{II}=$ 0.5, which corresponds to a binding energy of $\varepsilon_{I}^{(0)}/E_{n}=-0.5$. The spectral function of the impurity-impurity scattering matrix is shown in Fig. 2 as a function of $1/k_{n}a_{IB}$, which illustrates that the dressed dimer evolves from a well- to an ill-defined state. For weak boson-impurity interactions the spectral function is a sharp function with most of the spectral weight concentrated around the pole, giving a well-defined dressed dimer. As the polaron effects become more pronounced, we observe that the pole broadens significantly, hindering any signature of a dressed dimer state. This indicates that two-body bound states are well-defined when polaronic effects are weaker, whereas as polaron dressing increases, it prevents the emergence of well-defined dimers. In our numerical calculations, we add an imaginary term $i\gamma_{X}/E_{n}=0.1$ to the impurity Green’s function and use an energy cut-off of $\Lambda/E_{n}=144$ in the BSE.
@@
Our study relies on two approximations: First, the NSCT approximation, which incorporates the impurity-boson Feshbach resonance and is limited to a single Bogoliubov excitation out of the condensate. This formalism Rath and Schmidt (2013) has proven highly effective in describing Bose polaron experiments Jørgensen et al. (2016); Hu et al. (2016); Yan et al. (2020); Skou et al. (2021a), found to be in agreement with quantum Monte Carlo calculations Ardila et al. (2019), and has been widely used to study 2D Bose and Fermi polarons in semiconductors Sidler et al. (2017); Takemura et al. (2014); Levinsen et al. (2019); Bastarrachea-Magnani et al. (2019). Second, we neglect induced interactions between impurities, which can lead to bipolaron formation Naidon (2018); Camacho-Guardian et al. (2018); Camacho-Guardian and Bruun (2018). This approximation holds if the timescale for dimer formation via direct interactions is much shorter than that associated with mediated interactions. Specifically, this remains valid as long as the speed of sound in the BEC, $c_{s}=\sqrt{4\pi a_{BB}n_{0}/m_{b}^{2}},$ is much smaller than the typical velocity of the dimer, i.e., $1\gg\frac{c_{s}}{\sqrt{2|\varepsilon_{I}^{(0)}|/m_{b}}}\sim\sqrt{n_{0}a_{BB}a_{IB}^{2}}.$ Here, $a_{BB}$ is the boson-boson scattering length. We emphasize that, unlike mediated interactions, which are significantly affected by retardation effects, the direct two-body interaction is instantaneous.
@@
Conclusions.- We have studied the polaron effects on the formation of a bound state due to the direct interaction between two impurities. By using a Green’s function formalism which allowed us to the treat the single polaron non-perturbatively and by addressing the two-body problem by solving the BSE for the interaction of two impurity atoms we have demonstrated a subtle interplay between polaron physics and an underlying two-body bound state.
@@
Our results show that for weakly bound dimers, polaron effects can be strong enough to completely break the formation of a dimer whereas tightly bound dimers are more robust towards polaron in the strongly interacting impurity-boson regime. Our theoretical proposal can be tested with the current state-of-the-art experimental capabilities to probe Feshbach molecules Chin et al. (2010). An interesting avenue is to explore beyond the NSCT approximation the competition between polaron-few-body states Astrakharchik et al. (2021) and the formation of dressed dimers.
The study of the impact of beyond Fröhlich polaron dressing on two-body molecular states and bipolaron formation is potentially relevant to the study of Hubbard-Holstein models Wellein et al. (1996); Bonc˘a et al. (2000); Bonča and Trugman (2001); Macridin et al. (2004) or organic semiconductors.
I End Matter
We can illustrate our formalism in a suitable diagrammatic representation as in Fig. 6. In Fig. 6(a) we show the Bethe-Salpeter equation for the impurity-impurity scattering, which consists of the repeated impurity-impurity scattering with the direct interaction $a_{IB},$ depicted by the double wavy orange line. To solve the BSE, we obtain the impurity Green’s function propagator invoking a non self-consistent T-matrix approximation to describe the dressing of the impurities with the particles of the bath.
In Fig. 6(b) we represent diagrammatically the Dyson’s equation for the impurity propagator (solid blue line), which can be written in terms of the bare impurity propagator (dashed blue line) and the self-energy (pink circle), given by
$G_{cc}(\mathbf{p},\omega)=G^{(0)}_{cc}(\mathbf{p},\omega)+G^{(0)}_{cc}(\mathbf{p},\omega)\Sigma_{cc}(\mathbf{p},\omega)G_{cc}(\mathbf{p},\omega),$ (7)
with
$G^{(0)}_{cc}(\mathbf{p},\omega)=\frac{1}{\omega-\epsilon_{\mathbf{p}}^{(c)}},$
the ideal Green’s function for the impurity.
The self-energy is calculated under the NSCT approximation, and is represented in Fig. 6(c), here the red dashed lines correspond to condensate bosons.
$\Sigma_{cc}(\mathbf{p},\omega)=n_{0}\mathcal{T}(\mathbf{p},\omega).$ (8)
Finally, in Fig. 6(d) we show the diagrammatic representation of the T-matrix, giving the ladder approximation for the boson-impurity scattering $a_{IB}$ which we illustrate by the single wavy green line. For $a_{BB}=0,$ that is an ideal BEC, the impurity-boson T-matrix acquires a simple form:
$\mathcal{T}(\mathbf{p},\omega)=\frac{2\pi a_{IB}}{m_{r}}\frac{1}{1+\frac{i2\pi a_{IB}m_{r}^{3/2}}{m_{r}\sqrt{2}\pi}\sqrt{\omega-\frac{k^{2}}{2M}}}.$ (9)
Figure 6: (a) In orange, the BSE equation for the impurity-impurity scattering, the bare impurity-impurity interaction is depicted by the double wavy orange line and the impurity Green’s function by the solid blue line. (b) The Dyson’s equation for impurity Green’s function, the dashed line gives the ideal impurity Green’s function and the pink circle represents the self-energy, which is calculated following the NSCT approximation as in (b). The red dashed line correspond to condensate bosons whereas the solid red lines depict the propagator of the majority bosons. (d) Ladder approximation for the impurity-boson scattering.
One should note the similarities between the BSE and the T-matrix for impurity-impurity and impurity-boson scattering, respectively. Both account for the two-body physics at different levels: in the former, it provides the exact solution to the two-body problem; in the latter, it captures the Feshbach physics between the impurity and majority bosons. The NSCT provides the impurity Green’s function accounting for a non-perturbative dressing of the impurity with the bosons. The NSCT gives the input to solve the BSE for the impurity-impurity interaction retaining the full complexity of the scattering problem.